# Sysformer: Safeguarding Frozen Large Language Models with Adaptive System Prompts

### v1

@@ Abstract
As large language models (LLMs) are deployed in safety-critical settings, it is essential to ensure that their responses comply with safety standards. Prior research has revealed that LLMs often fail to grasp the notion of safe behaviors, resulting in either unjustified refusals to harmless prompts or the generation of harmful content. While substantial efforts have been made to improve their robustness, existing defenses often rely on costly fine-tuning of model parameters or employ suboptimal heuristic techniques. In this work, we take a novel approach to safeguard LLMs by learning to adapt the system prompts in instruction-tuned LLMs. While LLMs are typically pre-trained to follow a fixed system prompt, we investigate the impact of tailoring the system prompt to each specific user input on the safety of the responses. To this end, we propose Sysformer, a transformer model that updates an initial system prompt to a more robust system prompt in the LLM input embedding space while attending to the user prompt. While keeping the LLM parameters frozen, the Sysformer is trained to refuse to respond to a set of harmful prompts while responding ideally to a set of safe ones. Through extensive experiments on $5$ LLMs from different families and $2$ recent benchmarks, we demonstrate that Sysformer can significantly enhance the robustness of LLMs, leading to upto $80\%$ gain in the refusal rate on harmful prompts while enhancing the compliance with the safe prompts by upto $90\%$. Results also generalize well to sophisticated jailbreaking attacks, making LLMs upto $100\%$ more robust against different attack strategies. We hope our findings lead to cheaper safeguarding of LLMs and motivate future investigations into designing variable system prompts.
@@ 1 Introduction
Unregulated advancement of large language models (LLMs) poses extreme societal risks, such as automated warfare, societal inequalities, and misinformation (Bengio et al., 2024; Shevlane et al., 2023; Anwar et al., 2024; Chen & Shu, 2024). These risks largely arise from pre-training on vast unfiltered and uncurated Web data. It is therefore essential to develop safeguards to prevent the generation of potentially harmful content without compromising the beneficial applications.
While tuning models to align their responses with curated human preferences has been effective in ensuring safety (Ouyang et al., 2022; Ji et al., 2023), these approaches remain costly, requiring extensive human annotations, significant computational resources, and repeated retraining to adapt to new types of misuse. Furthermore, deeper vulnerabilities have been exposed through sophisticated jailbreaking techniques such as adversarial injection (Schwinn et al., 2024; Zou et al., 2023) and iterative prompt modification (Zhu et al., 2024; Chao et al., 2023). Thus, LLMs have been taught to respond safely to a specifically curated set of harm-inducing prompts by updating their parameters using finetuning (Mazeika et al., 2024), representation engineering (Zou et al., 2024), and adversarial finetuning (Jain et al., 2023). However, major gaps still exist as these approaches often generalize in unpredictable ways (Anwar et al., 2024; Qi et al., 2023) and may lead to forgetting useful pre-trained knowledge and refusal on benign prompts (Zhang & Wu, 2024; Wei et al., 2024).
Safeguards have thus also been developed on top of frozen LLMs using heuristic techniques such as smoothening responses over multiple input perturbations (Robey et al., 2023), system prompt tuning (Zheng et al., 2024), and prompt filtering (Kumar et al., 2023; Liu et al., 2024b; Jain et al., 2023). However, these strategies are often impractical since they either incur high computational costs through additional LLM calls or employ heuristic-based techniques removing potentially useful portions of user-provided prompts and suboptimal safeguarding.
@@
To address these concerns, we present Sysformer, a transformer-based architecture to adaptively modify the system prompt based on the user prompt for safeguarding frozen LLMs. Inspired by the multi-modal literature, we model the relationship between the user prompt and the system prompt by treating the embeddings of the system prompt as one modality and the user prompt embedding as the other. Thus, we learn to project the user prompt embeddings onto the system prompt while keeping the LLM frozen to refuse harmful prompts and comply with safe prompts in an end-to-end manner. Comprehensive experiments on $5$ LLMs and $2$ benchmarks show substantial improvement in the refusal rate on harmful prompts while preserving the performance on safe prompts. We also show that Sysformer can boost the robustness of LLMs over more sophisticated jailbreaking attack strategies as well by augmenting a few such examples during the training. Finally, we provide a detailed sensitivity analysis of different hyperparameters, training settings, and embedding architectures.
 Mazeika et al. (2024),  Jain et al. (2023),  Zou et al. (2024)  Kumar et al. (2023), Robey et al. (2023), Zhang et al. (2024)  Liu et al. (2024b),  Jain et al. (2023)  Zheng et al. (2024) Sysformer (proposed)
@@
User prompt unchanged
@@ 2 Related Work
Defense mechanisms. In response to the exposed vulnerabilities to prompt perturbations, various defensive mechanisms have been proposed in the literature. Finetuning-based strategies involve careful curation of adversarial harm-inducing user prompts along with safe prompts which are then used to update the parameters (Mazeika et al., 2024; Jain et al., 2023). Circuit breakers engineer the representations of harmful generations to cluster together for easier identification and pre-emptive stopping to avoid harm (Zou et al., 2024). While these strategies update the parameters, tuning-free methods have also been proposed, including paraphrasing the user prompts (Jain et al., 2023), removing tokens to maximize information bottleneck (Liu et al., 2024b), iteratively checking-and-erasing (Kumar et al., 2023), smoothening responses over multiple perturbations of user prompts (Robey et al., 2023), and simply asking the LLM to repeat its response (Zhang et al., 2024). Filtering-based strategies have led to the development of harm classifiers such as LlamaGuard (Inan et al., 2023), which are employed in both evaluation and content filtering. However, these defensive strategies either increase the computational cost through multiple calls or lead to arbitrary and strict filtering of the user prompts. For more flexible defenses, the system prompt has been tuned to maximize the generation of safe responses (Zheng et al., 2024). Here, we instead learn to adapt the system prompt based on the user prompt, enabling more efficient and context-aware safeguarding.
Lightweight Model Adaptation. Decoding-time methods such as IPA (Lu et al., 2023), Proxy-tuning (Liu et al., 2024a), and Bbox-adapter (Sun et al., 2024) are proposed to guide the token sampling of frozen models using fine-tuned smaller models for cheaper domain adaptation and reasoning. Frozen pre-trained vision and language models have been combined in a modular fashion by using a few self and cross-attention layers to enable multimodal capabilities (Li et al., 2023). Similarly, pre-trained LLMs have also been used as sentence representation by converting causal attention to bidirectional attention (BehnamGhader et al., 2024). In this work, we build upon Q-former (Li et al., 2023) and LLM2Vec (BehnamGhader et al., 2024) architectures to boost safety in frozen LLMs by learning a transformer to update the system prompt based on the user prompt.
@@ 3 Background and Problem
Consider an autoregressive large language model (LLM) $\mathcal{M}$ with an ordered vocabulary $\mathcal{V}$ of tokens. The model is pre-trained to predict the next token $x_{n+1}$ given a sequence of tokens $x_{1:n}$. Each token $x_{i}\in\mathcal{V}$ is first represented with an embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$ as $\mathbf{E}[x_{j}]=\mathbf{E}_{i}$, such that token $x_{j}$ comes at the index $i$ in $\mathcal{V}$. Then, these are transformed to obtain $\mathbf{Z}\in\mathbb{R}^{d}$ that is used to predict the next token using a trained density function $\mathbf{p}_{\mathcal{M}}(x_{n+1}|x_{1},x_{2},\cdots,x_{n})=\text{softmax}(\mathbf{W}\mathbf{Z})\in\mathbb{R}^{|\mathcal{V}|}$, where $\mathbf{W}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is either trainable or the same as the input embedding matrix $\mathbf{E}$. We thus use $\mathcal{M}(x_{1:n})$ to denote autoregressive sampling of tokens given $x_{1:n}$ using the density function $\mathbf{p}_{\mathcal{M}}$.
Modern LLMs are instruction-tuned with a default input that is prepended with the user prompt $\mathcal{P}:=p_{1:n}$. This is often called the system prompt $\mathcal{S}$, denoted as $s_{1:k}$ (Touvron et al., 2023). This implies the prediction is made as $\mathcal{M}(s_{1:k}\oplus p_{1:n})$ instead of just $\mathcal{M}(p_{1:n})$, where $\oplus$ concatenates the tokens together (the special tokens to identify the system prompt are ignored here for simplicity). This enables the deployer to reinforce certain guidelines of the intended usage that the model should not forget regardless of the user input. For example, one can explicitly write safety guidelines for safe responses and knowledge date cuts for reduced hallucinations.
@@
Suppose we have a pre-trained LLM $\mathcal{M}$ with frozen parameters. Then, the objective is to ensure without any filtering that it responds appropriately to benign user prompts but refuses to comply with any request $\tilde{x}_{1:n}=\tilde{x}_{1},\tilde{x}_{2},\cdots,\tilde{x}_{n}$ whose expected response can cause harm.
@@ 4 Method
We assume a supervised classification scenario where we have a set of user prompts $\{\mathcal{P}_{j}\}_{j=1}^{N}$, labeled either harmful or safe based on whether the intended response is harmful or safe respectively. Since the LLM encodes the system prompt as row-concatenated token embeddings $\mathbf{E}[\mathcal{S}]=\bigoplus_{i=1}^{k}\mathbf{E}[s_{i}]$, we further simplify the problem of combinatorial search over tokens $\widehat{\mathcal{S}}(\mathcal{P})$ to the problem of searching in a continuous space of $\widehat{\mathbf{S}}:=\mathbf{E}[\widehat{\mathcal{S}}]=\bigoplus_{i=1}^{k}\mathbf{E}[\hat{s}_{i}]\in\mathbb{R}^{k\times d}$. Thus, we relax Proposition 1 and propose
Proposition 2.
Given an LLM $\mathcal{M}$ with fixed/frozen parameters and a system prompt $\mathcal{S}$, our objective is to find a continuous, supervised, and adaptive system prompt $\widehat{\mathbf{S}}$ in the input embedding space of the LLM such that (1) $\widehat{\mathbf{S}}$ is some function of the user and initial system prompt, (2) LLM $\mathcal{M}$ generates “safe” responses for $\widehat{\mathbf{S}}\oplus\mathbf{E}[\mathcal{P}_{j}]$ over a set of user prompts $j\in[1,N]$.
@@ 4.1 Architecture
$\begin{cases}\text{Sysformer}_{\bm{\Theta}}(\mathcal{S},\mathcal{P};\mathbf{E}):=\widehat{\mathbf{S}}(\mathcal{S},\mathcal{P})=\widehat{\mathbf{S}}^{(L)},\\ \widehat{\mathbf{S}}^{(l)}=\text{CrossAttention}(\text{SelfAttention}(\widehat{\mathbf{S}}^{(l-1)}),\mathbf{P}),\\ \mathbf{P}:=\mathbf{E}[\mathcal{P}]=\mathbf{E}[p_{1}]\oplus\mathbf{E}[p_{2}]\oplus\cdots\oplus\mathbf{E}[p_{n}],\\ \widehat{\mathbf{S}}^{(0)}:=\mathbf{S}=\mathbf{E}[\mathcal{S}]=\mathbf{E}[s_{1}]\oplus\mathbf{E}[s_{2}]\oplus\cdots\oplus\mathbf{E}[s_{k}],\\ \end{cases}$ (1)
@@ 4.2 Training
The Sysformer parameters $\bm{\Theta}$ need to be trained such that Proposition 2 is satisfied, i.e., the LLM $\mathcal{M}$ generates safe responses for $\text{Sysformer}_{\bm{\Theta}}(\mathcal{S},\mathcal{P};\mathbf{E})\oplus\mathbf{E}[\mathcal{P}]$ for all user prompts $\mathcal{P}$ given the model’s input embedding $\mathbf{E}$. Thus, we propose different loss functions designed to induce this effect during training assuming access to a set of labeled user prompts $(\mathcal{P}_{i},y_{i})\in\mathcal{D}$ where $y_{i}=1$ means $\mathcal{P}_{i}$ is harmful and $y_{i}=0$ means $\mathcal{P}_{i}$ is a safe prompt.
@@
Distinguishing harmful and safe prompts. Following prior works (Zheng et al., 2024), we also enforce that LLM’s hidden representations can be used to distinguish between a harmful and a safe prompt. This is motivated by the fact that the refusal in LLMs has been shown to lie in a single direction of the representation space (Arditi et al., 2024) and thus, the prompt representations can also be linearly separated and aligned with the refusal direction for safety training. Thus, we train a linear layer $\mathbf{w}^{\top}\mathbf{x}+\mathbf{b}$ on top of the LLM’s final layer representation of the final token to classify between harmful and safe prompts. To do this, we use a binary cross-entropy loss and minimize $\mathcal{L}_{class}\propto\sum_{(\mathcal{P},y)\in\mathcal{D}}y\log\hat{y}+(1-y)\log\sigma(1-\hat{y})$, where $\hat{y}=\sigma(\mathbf{w}^{\top}\mathbf{Z}(\widehat{\mathbf{S}}(\mathcal{S},\mathcal{P})\oplus\mathbf{E}[\mathcal{P}])+\mathbf{b})$ and $\sigma(\cdot)$ is the sigmoid function.
@@
We consider a weighted combination of these loss functions to train the Sysformer parameters while keeping the LLM parameters frozen. In other words, we minimize $\mathcal{L}=w_{ref}\mathcal{L}_{ref}+w_{compl}\mathcal{L}_{compl}+w_{class}\mathcal{L}_{class}+w_{recon}\mathcal{L}_{recon}$ using gradient descent. Furthermore, we use self compliance loss if selfsafe is True, and employ additional compliance if add is True. Note that $\mathcal{L}_{add}$ is minimized separately after a single batch over $\mathcal{D}$ is completed. Algorithm 1 (Appendix B) describes the algorithm and different settings in more detail.

### v2

@@ Abstract
As large language models (LLMs) are deployed in safety-critical settings, it is essential to ensure that their responses comply with safety standards. Prior research has revealed that LLMs often fail to grasp the notion of safe behaviors, resulting in either unjustified refusals to harmless prompts or the generation of harmful content. While substantial efforts have been made to improve their robustness, existing defenses often rely on costly fine-tuning of model parameters or employ suboptimal heuristic techniques. In this work, we take a novel approach to safeguard LLMs by learning to adapt the system prompts in instruction-tuned LLMs. While LLMs are typically pre-trained to follow a fixed system prompt, we investigate the impact of tailoring the system prompt to each specific user input on the safety of the responses. To this end, we propose Sysformer  [^1] We release the code at https://github.com/Ksartik/sysformer., a transformer model that updates an initial system prompt to a more robust system prompt in the LLM input embedding space while attending to the user prompt. While keeping the LLM parameters frozen, the Sysformer is trained to refuse to respond to a set of harmful prompts while responding ideally to a set of safe ones. Through extensive experiments on $5$ LLMs from different families and $2$ recent benchmarks, we demonstrate that Sysformer can significantly enhance the robustness of LLMs, leading to upto $80\%$ gain in the refusal rate on harmful prompts while enhancing the compliance with the safe prompts by upto $90\%$. Results also generalize well to sophisticated jailbreaking attacks, making LLMs upto $100\%$ more robust against different attack strategies. We hope our findings lead to cheaper safeguarding of LLMs and motivate future investigations into designing variable system prompts.
@@ 1 Introduction
Unregulated advancement of large language models (LLMs) poses extreme societal risks, such as automated warfare, societal inequalities, and misinformation (Bengio et al., 2024; Shevlane et al., 2023; Anwar et al., 2024; Chen and Shu, 2024). It is therefore essential to develop safeguards to prevent the generation of potentially harmful content without compromising the beneficial applications. Although fine-tuning LLMs (Zou et al., 2024; Mazeika et al., 2024) offers some promise for aligning models with safety objectives, its limitations are increasingly evident, as deeper vulnerabilities continue to surface through sophisticated jailbreaking techniques (Zou et al., 2023; Chao et al., 2023). In practice, fine-tuning does not scale well with model size, generalizes unpredictably (Anwar et al., 2024; Qi et al., 2023), risks erasing useful pre-trained knowledge (Zhang and Wu, 2024), and often leads to overrefusal (Wei et al., 2024).
This highlights the need for safeguarding methods that avoid updating pretrained parameters. Existing approaches, such as making additional LLM calls for smoothening generations (Robey et al., 2023; Kumar et al., 2023), prompt filtering (Liu et al., 2024b; Jain et al., 2023), or post-generation moderation  [^2] https://platform.openai.com/docs/guides/moderation, offer valuable protection but often incur extra inference cost or risk filtering out useful content. Trainable modular attachments provide a complementary alternative since, by operating at the input level, they impose minimal overhead and avoid the rigidity of filtering, while still safeguarding frozen LLMs. Current designs, however, remain constrained by non-adaptive mechanisms (Zheng et al., 2024). In practice, many deployers have resorted to manual system-prompt tuning to enforce safe behavior, but this is labor-intensive, consumes context length, and is vulnerable to leakage  [^3] https://github.com/jujumilk3/leaked-system-prompts. These limitations call for more adaptive safeguarding mechanisms—ones that integrate the strengths of existing approaches while leveraging efficient, modular input-level defenses.
@@
To fill these gaps, we present Sysformer, a fixed-depth modular transformer architecture that attaches at the input of any LLM and adaptively modifies the system prompt based on the user prompt. Inspired by the multi-modal literature, we learn to transform the system prompt based on the user prompt by treating them as separate modalities such that the LLM refuses to harmful prompts and complies with safe prompts. Comprehensive experiments on $5$ LLMs and $2$ benchmarks show substantial improvement the refusal gap by increasing the refusal rate on harmful prompts and reducing it on benign prompts. We also show that Sysformer can boost the robustness of LLMs over more sophisticated jailbreaking attack strategies as well by augmenting a few such examples during the training. Finally, we provide a detailed sensitivity analysis of different hyperparameters, training settings, and embedding architectures.
Fine-tuning Smoothening Filtering SysEmbedder Sysformer
 Mazeika et al. (2024),  Jain et al. (2023),  Zou et al. (2024),  Zhang et al. (2024a)  Kumar et al. (2023), Robey et al. (2023), Zhang et al. (2024b)  Liu et al. (2024b),  Jain et al. (2023)  Zheng et al. (2024) (proposed)
@@
Unchanged user prompt
@@ 2 Related Work
Fine-tuning defenses. Different defensive mechanisms have been proposed in the literature to combat the exposed vulnerabilities to prompt perturbations. Finetuning-based strategies involve careful curation of adversarial harm-inducing user prompts along with safe prompts, which are then used to update the parameters so that the LLM generates appropriate outputs (Mazeika et al., 2024; Jain et al., 2023). Representation engineering methods propose alternative loss functions that directly manipulate the internal activations instead of the generated outputs, enabling localized low-rank parameter updates (Zou et al., 2024; Zhang et al., 2024a). On the other hand, our contribution is complementary to these advancements as we present a novel modular architecture that can be desirably trained with any loss function of choice while keeping the LLM parameters unchanged. Our proposed uniform modular attachment to any frozen LLM enables seamless compatibility with any fine-tuning approach, which can then be applied directly on top of Sysformer.
Frozen LLM defenses. Tuning-free methods have also been proposed that involve including paraphrasing the user prompts (Jain et al., 2023), adding a warning message (Xie et al., 2023), using in-context examples of jailbreaking (Wei et al., 2023), removing tokens to maximize information bottleneck (Liu et al., 2024b), iteratively checking-and-erasing (Kumar et al., 2023), smoothening responses over multiple perturbations of user prompts (Robey et al., 2023), and simply asking the LLM to repeat its response (Zhang et al., 2024b). Filtering-based strategies have led to the development of harm classifiers such as LlamaGuard (Inan et al., 2023), which are employed in both evaluation and content filtering. However, these defensive strategies either increase the computational cost through multiple calls or lead to arbitrary and strict filtering of the user prompts. For more flexible defenses, a modular approach has been proposed by tuning a single system prompt embedding to maximize the generation of safe responses (Zheng et al., 2024). Here, we instead learn to adapt the system prompt based on the user prompt, enabling more efficient and context-aware safeguarding.
Frozen Model Adaptation. Decoding-time methods such as IPA (Lu et al., 2023), Proxy-tuning (Liu et al., 2024a), and Bbox-adapter (Sun et al., 2024) are proposed to guide the token sampling of frozen models using fine-tuned smaller models for cheaper domain adaptation and reasoning. Frozen pre-trained vision and language models have been combined in a modular fashion by using a few self and cross-attention layers to enable multimodal capabilities (Li et al., 2023). Similarly, pre-trained LLMs have been adapted to embed sentences by converting causal attention to bidirectional attention (BehnamGhader et al., 2024). In this work, we build upon these architectures to boost safety in frozen LLMs by learning a module to update the system prompt based on the user prompt.
@@
Safety Benchmarks. Curation of high-quality harmful and safe prompts along with representative metrics is critical to understand and evaluate our progress in achieving safety. Thus, various resources and datasets have been developed for a systematic and comprehensive evaluation of LLM safety approaches (Chao et al., 2024; Souly et al., 2024; Mazeika et al., 2024; 2023; Wei et al., 2024). While performance on some of these benchmarks have been found to be confounded with other capabilities of scale (Ren et al., 2024), we use them to show gains in a large variety of fairly smaller LLMs.
@@ 3 Background and Problem
Consider an autoregressive LLM $\mathcal{M}$ with an ordered vocabulary $\mathcal{V}$ of tokens that predicts the next token $x_{n+1}$ given a sequence of tokens $x_{1:n}$. Each token $x_{i}\in\mathcal{V}$ is first represented with an embedding matrix $\mathbf{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$ as $\mathbf{E}[x_{j}]=\mathbf{E}_{i}$, such that token $x_{j}$ comes at the index $i$ in $\mathcal{V}$. Then, these are transformed through multiple layers to obtain $\mathbf{Z}\in\mathbb{R}^{d}$ that predicts the next token as $\mathbf{p}_{\mathcal{M}}(x_{n+1}|x_{1},x_{2},\cdots,x_{n})=\text{softmax}(\mathbf{W}\mathbf{Z})\in\mathbb{R}^{|\mathcal{V}|}$ for some $\mathbf{W}\in\mathbb{R}^{|\mathcal{V}|\times d}$. We thus use $\mathcal{M}(x_{1:n})$ to denote this autoregressive sampling of tokens given $x_{1:n}$ using the density function $\mathbf{p}_{\mathcal{M}}$.
Modern LLMs are instruction-tuned with a default input that is prepended with the user prompt $\mathcal{P}:=p_{1:n}$. This is often called the system prompt $\mathcal{S}$, denoted as $s_{1:k}$ (Touvron et al., 2023). This implies the prediction is made as $\mathcal{M}(s_{1:k}\oplus p_{1:n})$ instead of just $\mathcal{M}(p_{1:n})$, where $\oplus$ concatenates the tokens together (the special tokens to identify the system prompt are ignored here for simplicity). This enables the deployer to give general guidelines that the model should always follow.
@@
Given an LLM $\mathcal{M}$, we want to ensure that it responds naturally to benign user prompts but refuses to respond to harm-inducing user prompts, such that the trained parameters remain frozen, and user prompts remain unfiltered.
@@ 4 Method
Since the LLM encodes the system prompt as row-concatenated token embeddings $\mathbf{E}[\mathcal{S}]=\bigoplus_{i=1}^{k}\mathbf{E}[s_{i}]$, we further simplify the problem of combinatorial search over tokens $\widehat{\mathcal{S}}(\mathcal{P})$ to the problem of searching in a continuous space of $\widehat{\mathbf{S}}:=\mathbf{E}[\widehat{\mathcal{S}}]=\bigoplus_{i=1}^{k}\mathbf{E}[\hat{s}_{i}]\in\mathbb{R}^{k\times d}$. Thus, we further relax the above assumption by finding a continuous, supervised, and adaptive system prompt $\widehat{\mathbf{S}}$ in the input embedding space instead of $\widehat{\mathcal{S}}(\mathcal{P})$ in the textual space.
@@ 4.1 Architecture
[…]
@@ 4.2 Training
The Sysformer parameters $\bm{\Theta}$ need to be trained such that the LLM $\mathcal{M}$ generates safe responses for $\text{Sysformer}_{\bm{\Theta}}(\mathcal{S},\mathcal{P};\mathbf{E})\oplus\mathbf{E}[\mathcal{P}]$ for all user prompts $\mathcal{P}$. Thus, we propose different loss functions designed to induce this effect during training assuming access to a set of labeled user prompts $(\mathcal{P}_{i},y_{i})\in\mathcal{D}$ where $y_{i}=1$ means $\mathcal{P}_{i}$ is harmful and $y_{i}=0$ means $\mathcal{P}_{i}$ is a safe prompt.
@@
[…]
@@
We consider a weighted combination of these loss functions to train the Sysformer parameters while keeping the LLM parameters frozen. In other words, we minimize $\mathcal{L}=w_{ref}\mathcal{L}_{ref}+w_{compl}\mathcal{L}_{compl}+w_{class}\mathcal{L}_{class}+w_{recon}\mathcal{L}_{recon}$ using gradient descent. Furthermore, we use self compliance loss if selfsafe is True otherwise use fixed compliance, while if add is True, then additional compliance is used. Note that $\mathcal{L}_{add}$ is minimized separately after a single batch over $\mathcal{D}$ is completed. Algorithm 1 (Appendix B) describes the algorithm and different settings in more detail.